OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

hartnn (hartnn):

start by dividing -3 on both sides, what do u get ?

OpenStudy (imstuck):

I don't think s/he is here anymore!

OpenStudy (anonymous):

im here hold on let me work it out really quick

OpenStudy (anonymous):

i got X=-1 and X=1

hartnn (hartnn):

|2x+6| =4 2x+6 =4 , 2x+6 = -4 x=-1 is correct how did u get 1 ?

OpenStudy (anonymous):

ohhh i forgot to put the negative sign in front of the four

hartnn (hartnn):

see if u get -1,-5